OpenStudy (anonymous):

Find the value of x such that f(x)=12

OpenStudy (anonymous):

What the equation?

OpenStudy (anonymous):

\[f(x)=-\frac{ 1 }{ 5 }x+2\]

OpenStudy (anonymous):

So, we know f(x)= 12, then we can say 12= -1/5x + 2. Move the 2 over: 10 = -1/5x negate both sides: -10 = 1/5x Then multiply both sides by 5 -50 = 5/5x = x so: x=-50 Make sense?

OpenStudy (anonymous):

3*4